when a flower shop raised the price of a floral arrangement from $20 to $28, the number of the arrangements sold decreased from 30 Units a week to 20 Units a week. What is the price elasticity of demand for the flowers in this case? How did the shop's total revenue change as a result of this?
1
Expert's answer
2020-04-03T10:02:15-0400

Price elasticity of demand (PED) is the percentage change in quantity demanded of a commodity due to certain percentage change in its price.

PED = %change in quantity demand/%change in price

Mid point of quantity = "(Q1 + Q2)\/2" = (30 + 20)/2 = 25

Mid point of price = "(P1 + P2)\/2" = (20 +028)/2 = 24

%change in quantity = "(Q2 - Q1)\/Mid point quantity" = (20 - 30)/25 = -0.4

%change in quantity = "(P2 - P1)\/Mid point price" = (28 - 20)/24 = 0.3

Price Elasticity of Demand = -0.4/0.3 = 1.3


Total Revenue (TR) = "Total Quantity * Price per unit"

Initial TR = "Q1 * P1" = 30 * 20 = 600

Present TR = "Q2 * P2" = 20 * 28 = 560

The total revenue of the firm dropped from $600 to $560